If the person and the cricket ball consist of a system, find the position and velocity of the center of mass of the system at time t=0.5 s


1
Expert's answer
2021-12-13T18:33:14-0500

The position of the center of mass after 0.5s can be calculated using the 2nd equation of motion.

"y=y_0+v_yt+\frac{1}{2}gt^2"

"v_y=0"

g=9.8m/s

"y_0=11.5m"

t=0.5

"y=11.5+\frac{1}{2}×(-9.8)×0.5^2=10.275m"


Velocity of the center of mass after 0.5sec is calculated using the 1st equation of motion:

"v_y=u_y+at"

"u_g=0"

"v_y=-9.8\u00d70.5= -4.9ms^{-1}"

The negative sign signifies that the directions of motion